# Low-metallicity massive single stars with rotation

### v2

@@ 2.2 New PoWR spectra: WN and WO stars (no WC)
As for emission-line intensities, the mid- and late-pMS models (both WO) predict L 2.131037 erg s-1 and 3.771037 erg s-1, respectively. Since clumping introduces a factor of 3 uncertainty (as shown in Fig. 6), our use of the late-stage value to represent the entire post-main sequence is justified with the caveat that winds are probably not unclumped (cf. Roy et al. 2025) and that the observational range we attempt to match (L 2.2–5.51037 erg s-1) itself includes such an error margin.
@@ 2.3 Population synthesis
Equipped with all these model spectra, we now revisit the question of photoionization in I Zw 18 by building a synthetic population of metal-poor massive stars. Since chemically-homogeneous stars (TWUIN or O/WN/WO) are a straightforward outcome of rotating stellar evolution at the metallicity of I Zw 18, these stars are automatically included.
@@
It has been shown in Paper I that a population of single metal-poor massive stars on the main sequence is able to explain I Zw 18’s He II ionizing flux (Paper I, Sect. 10.4). No other observable was investigated in that simple population calculation. Further assumptions were (i) a Salpeter-type initial mass function (IMF); (ii) on-going star-formation for the past 3 Myr with a rate of 0.1 M⊙ yr-1 (this rate was observed by Lebouteiller et al. 2013) resulting a 300 000 M⊙ total mass in the region; (iii) an upper mass limit set to 500 M⊙; (iv) a rather high fraction, 20% of stars rotating fast enough to evolve chemically homogeneously and therefore becoming TWUIN stars (instead of cool, non-ionizing supergiants); and (v) black body emission333The question was followed up in a thesis (Szécsi 2016), where the black body emission was corrected for the wind optical depth according to Langer (1989)..
@@
We revise these assumptions here. Most importantly, we include the full evolution, i.e. both the main- and the post-main-sequences of our stellar models. We still use a regular Salpeter IMF (Salpeter 1955) for a continuous star-formation episode lasting for 3 Myr forming 0.1 M⊙ yr-1, same as before. The upper mass limit is, however, set to a more realistic 150 M⊙ (as opposed to 500 M⊙; though note that even 200 M⊙ might be acceptable in light of Bestenlehner et al. 2020). The ratio of chemically-homogeneously evolving stars is also relaxed: 10% now (as opposed to 20%), which is better supported by the rotational velocity distribution measured in the Small Magellanic Cloud (1/5 Z⊙, Mokiem et al. 2006) where about 10-15% of massive stars rotate faster than what is needed for chemically homogeneous evolution at the ten times lower metallicity of I Zw 18. This means that, in the population we build here, 90% of the massive stars are in fact ‘normal’, slow-rotating stars, i.e. not chemically-homogeneous, but classical single stellar evolution leading to O/B stars and supergiants. No classical WR stars though, as at this metallicity the supergiants do not lose their complete envelopes in the rather weak stellar winds, as shown by Szécsi (2016); Szécsi & Wünsch (2019); Szécsi et al. (2022). These ‘normally’ evolving O/B stars are taken into account assuming black body emission: their contribution to the ionizing flux is only significant near the zero-age main sequence when they are the hottest, while they do not contribute to emission line luminosities, for even in their early lives they are less hot than their chemically-homogeneously evolving counterparts (as seen in Figure 1, right panel), thus they surely experience transparent winds and absorptions-type spectra. We include low-mass stars ( 9 M⊙) only as mass holders down to 0.5 M⊙, the same way as in Paper I.
@@
To account for the radiation field emitted by the chemically-homogeneously evolving TWUIN/WO stars, we use the SED predicted by our synthetic spectra from Paper II with nominal mass-loss rate and smooth wind (Fig. 2). This way there is no need to correct for wind optical depth. Such a correction is a common practice for evolutionary models, done in e.g. Groh et al. (2019), Schootemeijer & Langer (2018), and Szécsi (2016). They followed the estimates of Langer (1989), which took only electron-scattering opacity into account: a fast but simplistic method. Since our spectra were created with state-of-the-art physical treatment of all elements in PoWR, including the high ionization stages of iron group elements (albeit with reduced abundances corresponding to the low metallicity), all sources of opacity are taken into account.
@@
As the spectra in Paper II are only provided for three initial masses, we linearly interpolate between the predicted flux values to get a smooth distribution. This is shown in Figure 5. The most massive stars’ post-main-sequence phase (spectral type WO) provides the largest ionizing photon count. So even if such stars are rare according to a typical IMF, their contribution to the He II ionization can be significant. Yet, the population synthesis includes the contribution of all the hot O-stars, as explained above. It depends on the star-formation history (see Sect. 3) whether very massive stars would form and dominate over the population or not. In either case, the ratio of WR/O stars in our synthetic population is around 0.001, which is consistent with the lowest-Z observations (upper limits) of López-Sánchez & Esteban (2010a, their Fig. 5).
@@ 2.4 Luminosity in emission lines
We perform spectral synthesis for emission lines by computing the line luminosity from our spectra. Figure 6 shows an example: the flux around 1550 is plotted for the models where the C iv line is in emission (that is, post-main-sequence models with nominal mass-loss rate).
@@ 3.1 Ionizing He II emission and C iv line luminosity
Table 1 summarizes our results. From our fiducial population built in Sects. 2.3–2.4, we predict a total He II ionizing flux of Q 1.801050 photons s-1 and a total C IV 1550 Å emission line luminosity of L1550 3.711037 erg s-1. These are not far from what is measured in I Zw 18: Q 1.331050 photons s-1 and L 2.2–5.51037 erg s-1 (cf. Sect. 1)444The range 2.25.51037 erg s-1 can be derived from the data presented in Brown et al. (2002) by taking the reported L=2.41037 erg s-1 for their Fig.1c and L=3.21037 erg s-1 for their Fig.1d (reproduced here in Figs. 11 & 14) and their statement “The ratio of C IV to He II emission in Figure 1c is 2.3 [while it is] only 0.7 in Figure 1d”, from which one gets L 5.51037 erg s-1 and L 2.21037 erg s-1for their Figs.1c and d, respectively.. Translating our statistics-based predictions into number of stars, we get about one or two post-main-sequence chemically-homogeneously-evolving objects—that is, WO stars, cf. Sect. 2.1-2.2—of mass 90 M⊙ in the population. The contribution of these one or two very massive WR type stars is responsible for the C IV line luminosity, while the He II photon emission is coming from them and the rest of the hot-star population (including normally-evolving models).
@@
There is no need for very massive stars, however, since alternative star-formation histories are within reasonable margins of theoretical and observational uncertainties. For example, a continuous episode lasting for 10 Myr with a much more moderate rate of only 0.03 M⊙ yr-1 also results in a 300 000 M⊙ region; this population gives us Q 1.961050 photons s-1 and L1550 3.431037 erg s-1.We have lowered the top mass value Mtop of the IMF from 150 to 120 M⊙ for this run, because already this more moderate value gives us the right predictions. In this case, the main source of the UV carbon line is made up by six chemically-homogeneously evolving WO stars of mass 40 M⊙ and another one of mass 80 M⊙; while the He+-emission is again explained by the combined contribution of these and the rest of the hot-star population (including those classified as type O in Fig. 1). So instead of a couple of very massive WO stars, a handful of more-or-less average ones can do the same job.
@@
We can relax our assumptions even further, by applying a 0.02 M⊙ yr-1 rate for a 15 Myr long star-formation episode (again with Mtop 120 M⊙, cf. Table 1). This way our 20 M⊙ model’s post-main-sequence phase (which hardens the resulting ionizing emission considerably, cf. Sect. 3.2) would be captured (the model’s total lifetime is 12.3 Myr). In such a scenario, about twenty-three WO stars of 20–25 M⊙ and two of around 50 M⊙ are responsible for a predicted carbon emission of L1550 2.831037 erg s-1 and an ionizing flux of Q 2.191050 photons s-1. Sect. 3.2 shows that this third scenario is the most preferred in terms of the spectral hardness, a further observational constraint. Table 1 presents a fourth and a fifth scenarios, which are even more moderate in terms of Mtop (80 an 40 M⊙), but they fail to account for the observed C iv luminosity.
@@
As for correcting for extinction, Brown et al. (2002) reports that E(B-V) 0.06 for an SMC-type extinction law fits the data well. From this one obtains a factor 2.3 attenuation at 1500. The recent study of Berg et al. (2022) quotes E(B-V) 0.07530.007, which gives a higher correction (2.7). With these numbers, one can estimate L (515)1037 erg s-1, which is around the same or up to 3 times higher than the L range we attempt to match here (2.2–5.51037 erg s-1). However, as Fig. 6 shows, introducing wind-clumping into our models increases our predicted line luminosity significantly—up to 3 times higher values with an extreme clumping factor. As explained in Sect. 5.2 among caveats, wind-clumping is one of the most underconstrained parameters of massive-star atmospheres, and supposing that the winds are completely unclumped (i.e. smooth) is a rather conservative assumption (cf. also Roy et al. 2025).
@@
We conclude therefore that a population of metal-poor massive single star models—including those with fast rotation that evolve to be TWUIN stars—can successfully explain the puzzling properties of the very dwarf galaxy, I Zw 18, that they were created for. The emission is dominated by WO stars, with a total C iv emission line intensity around 231037 erg s-1.
3.2 Nebular line ratios vs. continuum photon ratios
@@
The number of ionizing photons in I Zw 18 was obtained from measuring the strength of the line He II 4686 Å which originates in the ionized interstellar nebula. Kehrig et al. (2015b) report a line luminosity of 1.121038 erg s-1 from which they derive Q using common assumptions about the interstellar medium’s properties. An important observable to investigate is how the ratio of He II photons to Lyman continuum photons relate to the observed line intensities of the corresponding, excited nebular lines, He II 4686 Å and H 4861 Å. Observationally, this hardness ratio is measured to be in the range of I(4686)/I(H ) 0.020.04 (Izotov & Thuan 1998; Lebouteiller et al. 2017, although Kehrig et al. 2015b report values as high as 0.08 in individual spaxels). From a theoretical population of massive stars, one can predict this ratio as I(4686)/I(H ) AQ(He II)/Q(H I), where A is a factor taken to be 1.74 for typical nebular conditions (Stasińska et al. 2015).
@@
Izotov et al. (1997) | optical | 37006800 Å | MMTO 6.5 m | Fig. 8 |
Thuan & Izotov (2005) | optical | 32005200 Å | MMTO 6.5 m | Fig. 9 |
Kehrig et al. (2015b) | optical | 44405200 Å | PMAS 3.5 m (integral field) | Fig. 10 |
Brown et al. (2002) | UV | 11601710 Å | HST STIS | Fig. 11 |
Lecavelier des Etangs et al. (2004) | UV | 9101185 Å | FUSE | Fig. 12 |
Heap et al. (2015) / Berg et al. (2022) | UV | 11601645 Å | HST COS | Fig. 13 |
In our fiducial population (containing not just WO stars but all hot stars, including O-type stars from ‘normal’ evolution, cf. Sect. 2.3), the predicted photon count is Q(H I) 7.021052 s-1 in the H I continuum (as reported in Table 1) and Q(He II) 1.801050 s-1 in the He II continuum. From these, we derive the hardness ratio to be 0.004 (Table 1), which is an order of magnitude below what is observed (0.02-0.04).
However, for the alternative populations, when lower-mass stars are captured by relaxing the length of the starburst (Sect. 3.1), the value rises to 0.011 and 0.017, approaching the expected range. From this we deduce that the extremely short and intense scenario suggested by Paper I (our fiducial population) where a couple of very massive stars dominate the emission is rendered unlikely. Instead, a population where all massive stars’ post-main-sequence phases are contributing, not just the youngest (very massive) ones, should be considered realistic. In our second scenario (SF-10 in Table 1) the age of the starburst excludes post-main-sequence objects with initial masses below 30 M⊙, while in our third scenario (SF-15 in Table 1), everything down to 20 M⊙ is incorporated. The contribution of this evolutionary model to the hardness of the population is demonstrated by Figure 7, where evolutionary sequences are compared in terms of their (black-body-estimated) ionizing emissions. We suggest that in the future, PoWR models down to 12 M⊙ could be computed and tested for their effect of increasing the hardness—an apparent trend—fitting the observed range even better.
Since lowering Mtop from 150 to 120 M⊙ helped us match the measured hardness, we checked even lower values, down to 80 and even 40 M⊙. They give less good matches: the general trend is that lowering Mtop does not influence our He ii and hardness predictions at all (at least as long as the 15 Myr star-formation history is applied), but the C iv line luminosity decreases below the observed range.
We conclude that our theory is consistent with the measured hardness in I Zw 18, and by matching this value as an additional constraint we can draw important conclusions on the length/intensity of the starbursts and the stellar content, favouring less violent and more relaxed scenarios over those dominated by very-massive stars.

### v3

@@ 2.2 New PoWR spectra: WN and WO stars (no WC)
As for emission-line intensities, the mid- and late-pMS models (both WO) predict L 2.131037 erg s-1 and 3.771037 erg s-1, respectively. Since clumping introduces a factor of 3 uncertainty (as is shown in Fig. 6), our use of the late-stage value to represent the entire post-main-sequence is justified with the caveat that winds are probably not unclumped (see Roy:2025) and that the observational range we attempt to match (L 2.2–5.51037 erg s-1) itself includes such an error margin.
@@ 2.3 Population synthesis
Equipped with all these model spectra, we now revisit the question of photoionization in I Zw 18 by building a synthetic population of metal-poor massive stars. Since chemically homogeneous stars (TWUIN or O/WN/WO) are a straightforward outcome of rotating stellar evolution at the metallicity of I Zw 18, these stars are automatically included.
@@
It was shown in Szecsi:2015 that a population of single metal-poor massive stars on the main sequence is able to explain I Zw 18’s He II ionizing flux (Szecsi:2015, Sect. 10.4). No other observable was investigated in that simple population calculation. Further assumptions were (i) a Salpeter-type initial mass function (IMF); (ii) ongoing star formation for the past 3 Myr with a rate of 0.1 M⊙ yr-1 (this rate was observed by Lebouteiller:2013) resulting a 300 000 M⊙ total mass in the region; (iii) an upper mass limit set to 500 M⊙; (iv) a rather high fraction, 20% of stars rotating fast enough to evolve chemically homogeneously and therefore becoming TWUIN stars (instead of cool, non-ionizing supergiants); and (v) black body emission444The question was followed up in a thesis (Szecsi:2016), where the black body emission was corrected for the wind optical depth according to Langer:1989a..
@@
We revise these assumptions here. Most importantly, we included the full evolution, i.e. both the main- and the post-main-sequences of our stellar models. We still used a regular Salpeter IMF (Salpeter:1955) for a continuous star formation episode lasting for 3 Myr forming 0.1 M⊙ yr-1, same as before. The upper mass limit was, however, set to a more realistic 150 M⊙ (as opposed to 500 M⊙; though note that even 200 M⊙ might be acceptable in light of Bestenlehner:2020). The ratio of chemically homogeneously evolving stars was also relaxed: 10% now (as opposed to 20%), which is better supported by the rotational velocity distribution measured in the Small Magellanic Cloud (1/5 Z⊙, Mokiem:2006), where about 10-15% of massive stars rotate faster than what is needed for CHE at the ten times lower metallicity of I Zw 18. This means that, in the population we built here, 90% of the massive stars are in fact ‘normal’, slow-rotating stars, i.e. not chemically homogeneous, but classical single stellar evolution leading to O/B stars and supergiants. No classical WR stars though, as at this metallicity the supergiants do not lose their complete envelopes in the rather weak stellar winds, as is shown by Szecsi:2016; Szecsi:2019; Szecsi:2022ok. These ‘normally’ evolving O/B stars are taken into account assuming black body emission: their contribution to the ionizing flux is only significant near the zero-age main sequence (ZAMS) when they are the hottest, while they do not contribute to emission line luminosities, for even in their early lives they are less hot than their chemically homogeneously evolving counterparts (as is seen in Figure 1, right panel); thus, they surely experience transparent winds and absorption-type spectra. We include low-mass stars ( 9 M⊙) only as mass holders down to 0.5 M⊙, the same way as in Szecsi:2015.
@@
To account for the radiation field emitted by the chemically homogeneously evolving TWUIN/WO stars, we used the SED predicted by our synthetic spectra from Kubatova:2019 with a nominal mass-loss rate and smooth wind (Fig. 2). This way there is no need to correct for wind optical depth. Such a correction is a common practice for evolutionary models, done in for example Groh:2019, Schootemeijer:2018, and Szecsi:2016. They followed the estimates of Langer:1989a, which took only electron-scattering opacity into account: a fast but simplistic method. Since our spectra were created with state-of-the-art physical treatment of all elements in PoWR, including the high ionization stages of iron group elements (albeit with reduced abundances corresponding to the low metallicity), all sources of opacity are taken into account.
@@
As the spectra in Kubatova:2019 are only provided for three initial masses, we linearly interpolated between the predicted flux values to get a smooth distribution. This is shown in Figure 5. The most massive stars’ post-main-sequence phase (spectral type WO) provides the largest ionizing photon count. So even if such stars are rare according to a typical IMF, their contribution to the He II ionization can be significant. Yet, the population synthesis includes the contribution of all the hot O-stars, as explained above. It depends on the star formation history (see Sect. 3) whether very massive stars would form and dominate over the population or not. In either case, the ratio of WR/O stars in our synthetic population is around 0.001, which is consistent with the lowest-Z observations (upper limits) of LopezSanchez:2010a.
To sum up, all our assumptions about the synthetic population are either less extreme than those in Szecsi:2015 or identical, while the models we apply are more elaborate. Our results, including further parameter studies, are reported and compared to observational data in Sect. 3.
@@ 2.4 Luminosity in emission lines
We performed spectral synthesis for emission lines by computing the line luminosity from our spectra. Figure 6 shows an example: the flux around 1550 is plotted for the models where the C iv line is in emission (that is, post-main-sequence models with a nominal mass-loss rate).
@@ 3.1 Ionizing He II emission and C iv line luminosity
Table 1 summarizes our results. From our fiducial population built in Sects. 2.3–2.4, we predict a total He II ionizing flux of Q 1.801050 photons s-1 and a total C IV 1550 Å emission line luminosity of L1550 3.711037 erg s-1. These are not far from what is measured in I Zw 18: Q 1.331050 photons s-1 and L 2.2–5.51037 erg s-1 (see Sect. 1)555The range 2.25.51037 erg s-1 can be derived from the data presented in Brown:2002 by taking the reported L=2.41037 erg s-1 for their Fig.1c and L=3.21037 erg s-1 for their Fig.1d (reproduced here in Figs. 11 & 14) and their statement “The ratio of C IV to He II emission in Figure 1c is 2.3 [while it is] only 0.7 in Figure 1d”, from which one gets L 5.51037 erg s-1 and L 2.21037 erg s-1for their Figs.1c and d, respectively.. Translating our statistics-based predictions into number of stars, we get about one or two post-main-sequence chemically homogeneously evolving objects – that is, WO stars (see Sect. 2.1-2.2) – of mass 90 M⊙ in the population. The contribution of these one or two very massive WR type stars is responsible for the C IV line luminosity, while the He II photon emission is coming from them and the rest of the hot-star population (including normally evolving models).
@@
There is no need for very massive stars, however, since alternative star formation histories are within reasonable margins of theoretical and observational uncertainties. For example, a continuous episode lasting for 10 Myr with a much more moderate rate of only 0.03 M⊙ yr-1 also results in a 300 000 M⊙ region; this population gives us Q 1.961050 photons s-1 and L1550 3.431037 erg s-1.We have lowered the top mass value Mtop of the IMF from 150 to 120 M⊙ for this run, because already this more moderate value gives us the right predictions. In this case, the main source of the UV carbon line is made up by six chemically homogeneously evolving WO stars of mass 40 M⊙ and another one of mass 80 M⊙; while the He+-emission is again explained by the combined contribution of these and the rest of the hot-star population (including those classified as type O in Fig. 1). So instead of a couple of very massive WO stars, a select number of more or less average ones can do the same job.
@@
We can relax our assumptions even further, by applying a 0.02 M⊙ yr-1 rate for a 15 Myr long star formation episode (again with Mtop 120 M⊙; see Table 1). This way our 20 M⊙ model’s post-main-sequence phase (which hardens the resulting ionizing emission considerably; see Sect. 3.2) would be captured (the model’s total lifetime is 12.3 Myr). In such a scenario, about twenty-three WO stars of 20–25 M⊙ and two of around 50 M⊙ are responsible for a predicted carbon emission of L1550 2.831037 erg s-1 and an ionizing flux of Q 2.191050 photons s-1. Sect. 3.2 shows that this third scenario is the most preferred in terms of the spectral hardness, a further observational constraint. Table 1 presents a fourth and a fifth scenarios, which are even more moderate in terms of Mtop (80 an 40 M⊙), but they fail to account for the observed C iv luminosity.
@@
As for correcting for extinction, Brown:2002 reported that E(B-V) 0.06 for an SMC-type extinction law fits the data well. From this one obtains a factor 2.3 attenuation at 1500. The recent study of Berg:2022 quotes E(B-V) 0.07530.007, which gives a higher correction (2.7). With these numbers, one can estimate L (515)1037 erg s-1, which is around the same or up to 3 times higher than the L range we attempt to match here (2.2–5.51037 erg s-1). However, as Fig. 6 shows, introducing wind-clumping into our models increases our predicted line luminosity significantly – up to 3 times higher values with an extreme clumping factor. As is explained in Sect. 5.2 among caveats, wind-clumping is one of the most underconstrained parameters of massive-star atmospheres, and supposing that the winds are completely unclumped (i.e. smooth) is a rather conservative assumption (see also Roy:2025).
@@
We conclude therefore that a population of metal-poor massive single star models – including those with fast rotation that evolve to be TWUIN stars – can successfully explain the puzzling properties of the very dwarf galaxy, I Zw 18, which they were created for. The emission is dominated by WO stars, with a total C iv emission line intensity around 231037 erg s-1.
3.2 Nebular line ratios versus continuum photon ratios
@@
The number of ionizing photons in I Zw 18 was obtained from measuring the strength of the line He II 4686 Å which originates in the ionized interstellar nebula. Kehrig:2015 report a line luminosity of 1.121038 erg s-1 from which they derive Q using common assumptions about the interstellar medium’s properties. An important observable to investigate is how the ratio of He II photons to Lyman continuum photons relate to the observed line intensities of the corresponding, excited nebular lines, He II 4686 Å and H 4861 Å. Observationally, this hardness ratio is measured to be in the range of I(4686)/I(H ) 0.020.04 (Izotov:1998; Lebouteiller:2017, although Kehrig:2015 report values as high as 0.08 in individual spaxels). From a theoretical population of massive stars, one can predict this ratio as I(4686)/I(H ) AQ(He II)/Q(H I), where A is a factor taken to be 1.74 for typical nebular conditions (Stasinska:2015).
@@
| Izotov:1997 | optical | 37006800 Å | MMTO 6.5 m | Fig. 8 |
| Thuan:2005 | optical | 32005200 Å | MMTO 6.5 m | Fig. 9 |
| Kehrig:2015 | optical | 44405200 Å | PMAS 3.5 m (integral field) | Fig. 10 |
| Brown:2002 | UV | 11601710 Å | HST STIS | Fig. 11 |
| LecavelierdesEtangs:2004 | UV | 9101185 Å | FUSE | Fig. 12 |
| Heap:2015 / Berg:2022 | UV | 11601645 Å | HST COS | Fig. 13 |
In our fiducial population (containing not just WO stars but all hot stars, including O-type stars from ‘normal’ evolution; see Sect. 2.3), the predicted photon count is Q(H I) 7.021052 s-1 in the H I continuum (as reported in Table 1) and Q(He II) 1.801050 s-1 in the He II continuum. From these, we derive the hardness ratio to be 0.004 (Table 1), which is an order of magnitude below what is observed (0.02-0.04).
However, for the alternative populations, when lower-mass stars are captured by relaxing the length of the starburst (Sect. 3.1), the value rises to 0.011 and 0.017, approaching the expected range. From this we deduce that the extremely short and intense scenario suggested by Szecsi:2015 (our fiducial population) where a couple of very massive stars dominate the emission is rendered unlikely. Instead, a population where all massive stars’ post-main-sequence phases are contributing, not just the youngest (very massive) ones, should be considered realistic. In our second scenario (SF-10 in Table 1) the age of the starburst excludes post-main-sequence objects with initial masses below 30 M⊙, while in our third scenario (SF-15 in Table 1), everything down to 20 M⊙ is incorporated. The contribution of this evolutionary model to the hardness of the population is demonstrated by Figure 7, where evolutionary sequences are compared in terms of their (black-body-estimated) ionizing emissions. We suggest that in the future, PoWR models down to 12 M⊙ could be computed and tested for their effect of increasing the hardness – an apparent trend – fitting the observed range even better.
Since lowering Mtop from 150 to 120 M⊙ helped us match the measured hardness, we checked even lower values, down to 80 and even 40 M⊙. They give less good matches: the general trend is that lowering Mtop does not influence our He ii and hardness predictions at all (at least as long as the 15 Myr star formation history is applied), but the C iv line luminosity decreases below the observed range.
We conclude that our theory is consistent with the measured hardness in I Zw 18, and by matching this value as an additional constraint we can draw important conclusions on the length/intensity of the starbursts, as well as the stellar content, favouring less violent and more relaxed scenarios over those dominated by very-massive stars.